If P(X) = .23 and P(X and Y) = .12 and P(X or Y) =.34, find the complement of P(Y)?
Answer by ikleyn(52780)

.

You are given


    P(X) = 0.23

    P(X and Y) = 0.12

    P(X or Y) = 0.34 = P(X) + P(Y) - P(X and Y) = 0.23 + P(Y) - 0.12.


From the last equation,  


    P(Y) = 0.34 - 0.23 + 0.12 = 0.23.


Then the complement of P(Y) is


    P'(Y) = 1 - P(Y) = 1 - 0.23 = 0.77.    ANSWER

Solved.